OpenStudy (anonymous):

The rectangular floor of a storage building had a length of 25 yards and a perimeter of 90 yards. A construction crew knocked out one wall and extended the length of the floor by 10 yards. What was the area of the new floor?

OpenStudy (anonymous):

well perimeter P is 2x(L+W) where L is length and W is width L is 25, P is 90 = you can figure out W easily from there. then use found W and L+10 to find out area

OpenStudy (anonymous):

Yeah, but I'm confused so what seems easy to you may not seem easy to me... but thanks anyway for taking the time to answer!

OpenStudy (anonymous):

the purpose for this site is to give help, not answers, otherwise this wont be a learning site

OpenStudy (anonymous):

if you are confused - ask ill try to hint more

OpenStudy (anonymous):

yes I understand, but I don't understand how to find the width. I'm new to this subject, so I was hoping for a more descriptive answer.

OpenStudy (anonymous):

Sorry, I don't mean to be rude thanks for taking the time to help me out though :)

OpenStudy (anonymous):

finding W is same as finding x: P=2(L+W) can be easily written as 90=2(25+x) and then solved for x

OpenStudy (anonymous):

(sorry I don't mean to sound stupid, i want to keep my straight As lol!) but how do I solve for x...?

OpenStudy (anonymous):

you need to isolate x on one side to obtain the answer well here is basic example 5x-7=14 first you add 7 to both sides (5x-7)+7 = (14)+7 5x = 21 then you divide both sides by 5 5x/5 = 21/5 x = 21/5 There you go you just solved for x and found out, that x is equals to 21/5

OpenStudy (anonymous):

oh I see . Thanks so much for the help! I gave you best answer :) And I fanned you!

OpenStudy (anonymous):

good luck - remember if you want straight As dont seek just answer here - this site wont be with you on the test =p Its much better to learn how to solve any problem like this, than to know answer for single question.